OpenStudy (anonymous):

What is the solution to r/4 = 1/3?

OpenStudy (sasogeek):

\[\large \frac{r}{4}=\frac{1}{3} \ \ \ \ \ \ \ multiply \ through \ by \ 4 \]\[\large 4(\frac{r}{4})=4(\frac {1}{3}) \ \ \ \ \ 4 \ will \ cancel \ out \]\[\large r=\frac{4}{3} \]